# Crossed products by compact group actions with the weak tracial Rokhlin property

### v2

@@ 6. The weak tracial Rokhlin property
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the naive weak tracial Rokhlin property if for any $\varepsilon>0$, any finite set $F\subseteq A$, any finite set $S\subseteq C(G)$, and any $x\in A_{+}$ with $\|x\|=1$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
Similar to condition (4) in [43, Definition 4.1], since we consider the transfer of properties from $A$ to $A^{\alpha}$, the Cuntz comparison condition in $A^{\alpha}$ is necessary to handle certain technical details. Thus, the condition $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$ appears essential; without it, we may not be able to complete most of the previous proofs.
@@
Let $A$ be a unital simple infinite-dimensional C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be a finite group action with the weak tracial Rokhlin property. Then for any finite set $F\subseteq A$, any $\varepsilon>0$, any $x\in A_{+}$ with $\|x\|=1$, and any $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and contractive positive elements $(d_{g})_{g\in G}$ such that
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [43, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$. Define continuous functions $h,h_{0}\colon[0,1]\rightarrow[0,1]$ by
@@
Then we have $\|x-h_{0}(x)\|\leq\frac{\varepsilon}{2}$. Moreover, since $\|x\|=1$, we have $h(x)\neq 0$. By [43, Lemma 2.6], we can choose a nonzero positive element $x_{0}\in\overline{h(x)Ah(x)}$ such that $x_{0}\precsim_{A}z$. We may assume $\|x_{0}\|=1$.
Apply Lemma 3.2 to $\alpha$ for given $x_{0}$, $\frac{\varepsilon}{2}$, and $F$. We obtain contractive positive elements $(d_{g})_{g\in G}\in A$ and $d\in A^{\alpha}$ such that
@@
By [7, Lemma 3.7], $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}z$. Since $z\in\overline{yA^{\alpha}y}$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$, which is condition (6).
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [43, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$.
@@
Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist $n+1$ mutually orthogonal, Cuntz equivalent nonzero positive elements in $A^{\alpha}$. Let $e$ be one of them. Thus, by [43, Lemma 2.1], there exists $r\in(\overline{eA^{\alpha}e})_{+}\setminus\{0\}$ such that 0 is a limit point of $\mathrm{sp}(r)$. Define a continuous function $h\colon[0,1]\to[0,1]$ by
@@
Apply [43, Lemma 2.6] to find $x_{0}\in\overline{h(x)Ah(x)}$ such that
@@
Apply Lemma 3.2 to $\alpha$ for given $x_{0}$, $\frac{\varepsilon}{2}$, and $F$. We obtain contractive positive elements $(d_{g})_{g\in G}\in A$ and $d\in A^{\alpha}$ such that
@@
By Proposition 6.3, we only need to show that $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. Since $(1-d-\varepsilon)_{+}\precsim_{A}x_{0}\precsim_{A}r$, by [7, Lemma 3.7], we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}r$. Thus $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}e$. Let $\tau\in\mathrm{QT}(A^{\alpha})$. Then $d_{\tau}((1-d-\varepsilon)_{+})\leq d_{\tau}(e)$, and by Lemma 2.7,
@@
Using [7, Theorem 4.1] in the first step and (6.1) in the second step, we have $\mathrm{rc}(A^{\alpha})\leq\mathrm{rc}(A)<1-\frac{2}{n}$. Since (6.2) holds for all $\tau\in\mathrm{QT}(A^{\alpha})$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the weak tracial Rokhlin property if for any $\varepsilon>0$, any finite set $F\subseteq A$, any finite set $S\subseteq C(G)$, any $x\in A_{+}$ with $\|x\|=1$, and any $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a contractive positive element $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@ 7. A totally disconnected compact group action
By [34, Theorem 1.10.16], $A\cong\mathcal{Z}$.
Definition 7.3.
@@
Then for every $\varepsilon>0$, every finite set $F\subseteq A$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal contractive positive elements $(d_{g})_{g\in G}\in A$ satisfying $d=\sum_{g\in G}d_{g}\in A^{\alpha}$, such that
@@
This follows from [28, Example 5.10] (see also [1, Theorem 3.10]) and Proposition 6.3. ∎
@@
Let $n\in\mathbb{N}$, let $A_{1},A_{2},\dots,A_{n}$ be unital C*-algebras, and for $m=1,2,\dots,n$ let $e_{m}\in A_{m}$ be a contractive positive element and let $\tau_{m}$ be a tracial state on $A_{m}$. Let $A=A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}$ (minimal tensor product), and let
@@
Let $n\in\mathbb{N}$, let $A_{1},A_{2},\dots,A_{n}$ be unital C*-algebras, and for $m=1,2,\dots,n$ let $e_{m}\in A_{m}$ be a contractive positive element and let $\tau_{m}$ be a tracial state on $A_{m}$. Let $A=A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}$ (minimal tensor product), and let
@@
Let $F\subseteq A$ and $S\subseteq C(G)$ be finite sets, let $\varepsilon>0$, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Without loss of generality, we may assume $\|a\|\leq 1$ for all $a\in F$, $\|f\|\leq 1$ for all $f\in S$, and $\varepsilon<1$. We need to find a contractive positive element $d\in A^{\alpha}$ and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that
@@
By [28, Example 5.10], the action $\gamma$ has the weak tracial Rokhlin property. Therefore, by [13, Lemma 3.3], $B\rtimes_{\gamma}S_{2}$ is simple, and hence by Theorem 4.5, $B^{\gamma}$ is simple. It is easy to verify that $A^{\alpha}$ can be identified with $\bigotimes_{m=1}^{\infty}B_{m}^{\gamma}$ (see [38, Lemma 4.5]), hence it is simple. Meanwhile, since $B$ has strict comparison, $B^{\gamma}$ also has it, and thus $A^{\alpha}$ does as well (see [12, Proposition 5.16.1(4)]). Since $A^{\alpha}$ is unital, stably finite, and amenable, $\mathrm{T}(A^{\alpha})$ is a nonempty compact set. We denote $\delta_{2}=\inf_{\tau\in\mathrm{T}(A^{\alpha})}d_{\tau}(y)>0$.
@@
Since the action $\gamma$ has the weak tracial Rokhlin property, by Lemma 7.4, for each $1\leq i\leq N$, set $F_{i}=\{c_{ij}\colon 1\leq j\leq k\}$, we obtain a contractive positive element $d^{(i)}\in B^{\gamma}$ and mutually orthogonal contractive positive elements $(d_{g}^{(i)})_{g\in H_{i}}\in B$ such that
@@
In fact, we can use the above technique to construct examples of permutation actions on unital simple tracially $\mathcal{Z}$-stable C*-algebras with few projections. For example, the algebras constructed in [14] are all unital simple tracially $\mathcal{Z}$-stable with few projections, have strict comparison, and stable rank one (hence stably finite).

### v3

@@ 6. The weak tracial Rokhlin property
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the naive weak tracial Rokhlin property if for any $\varepsilon>0$, any finite set $F\subseteq A$, any finite set $S\subseteq C(G)$, and any $x\in A_{+}$ with $\|x\|=1$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
Similar to condition (4) in [42, Definition 4.1], since we consider the transfer of properties from $A$ to $A^{\alpha}$, the Cuntz comparison condition in $A^{\alpha}$ is necessary to handle certain technical details. Thus, the condition $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$ appears essential; without it, we may not be able to complete most of the previous proofs.
@@
Let $A$ be a unital simple infinite-dimensional C*-algebra. Let $\alpha\colon G\to\mathrm{Aut}(A)$ be a finite group action with the weak tracial Rokhlin property. Then for any finite set $F\subseteq A$, any $\varepsilon>0$, any $x\in A_{+}$ with $\|x\|=1$, and any $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and positive contractions $(d_{g})_{g\in G}$ such that the following hold:
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [42, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$. Define continuous functions $h,h_{0}\colon[0,1]\rightarrow[0,1]$ by
@@
Then we have $\|x-h_{0}(x)\|\leq\frac{\varepsilon}{2}$. Moreover, since $\|x\|=1$, we have $h(x)\neq 0$. By [42, Lemma 2.6], we can choose a nonzero positive element $x_{0}\in\overline{h(x)Ah(x)}$ such that $x_{0}\precsim_{A}z$. We may assume $\|x_{0}\|=1$.
Apply Lemma 3.2 to $\alpha$ for given $x_{0}$, $\frac{\varepsilon}{2}$, and $F$. We obtain positive contractions $(d_{g})_{g\in G}\in A$ and $d\in A^{\alpha}$ such that the following hold:
@@
By [6, Lemma 3.7], $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}z$. Since $z\in\overline{yA^{\alpha}y}$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}y$, which is condition (6).
@@
Let $\varepsilon>0$, let $F\subseteq A$ be a finite set, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Since $A^{\alpha}$ is simple and non-type I, by [42, Lemma 2.1], there exists a positive element $z\in\overline{yA^{\alpha}y}$ such that 0 is a limit point of $\mathrm{sp}(z)$.
@@
Since $A^{\alpha}$ is simple and non-type I, by Lemma 2.6, there exist $n+1$ mutually orthogonal, Cuntz equivalent nonzero positive elements in $A^{\alpha}$. Let $e$ be one of them. Thus, by [42, Lemma 2.1], there exists $r\in(\overline{eA^{\alpha}e})_{+}\setminus\{0\}$ such that 0 is a limit point of $\mathrm{sp}(r)$. Define a continuous function $h\colon[0,1]\to[0,1]$ by
@@
Apply [42, Lemma 2.6] to find $x_{0}\in\overline{h(x)Ah(x)}$ such that
@@
Apply Lemma 3.2 to $\alpha$ for given $x_{0}$, $\frac{\varepsilon}{2}$, and $F$. We obtain positive contractions $(d_{g})_{g\in G}\in A$ and $d\in A^{\alpha}$ such that the following hold:
@@
By Proposition 6.3, we only need to show that $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. Since $(1-d-\varepsilon)_{+}\precsim_{A}x_{0}\precsim_{A}r$, by [6, Lemma 3.7], we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}r$. Thus $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}e$. Let $\tau\in\mathrm{QT}(A^{\alpha})$. Then $d_{\tau}((1-d-\varepsilon)_{+})\leq d_{\tau}(e)$, and by Lemma 2.7,
@@
Using [6, Theorem 4.1] in the first step and (6.1) in the second step, we have $\mathrm{rc}(A^{\alpha})\leq\mathrm{rc}(A)<1-\frac{2}{n}$. Since (6.2) holds for all $\tau\in\mathrm{QT}(A^{\alpha})$, we have $(1-d-\varepsilon)_{+}\precsim_{A^{\alpha}}d$. ∎
@@
Let $G$ be a second-countable compact group, let $A$ be a unital simple infinite-dimensional C*-algebra, and let $\alpha\colon G\to\mathrm{Aut}(A)$ be an action. We say that the action $\alpha$ has the weak tracial Rokhlin property if for any $\varepsilon>0$, any finite set $F\subseteq A$, any finite set $S\subseteq C(G)$, any $x\in A_{+}$ with $\|x\|=1$, and any $y\in(A^{\alpha})_{+}\setminus\{0\}$, there exist a positive contraction $d\in A^{\alpha}$ and a contractive completely positive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@ 7. A totally disconnected compact group action
By [33, Theorem 1.10.16], $A\cong\mathcal{Z}$.
Notation 7.3.
@@
Then for every $\varepsilon>0$, every finite set $F\subseteq A$, and every $x\in A_{+}$ with $\|x\|=1$, there exist mutually orthogonal positive contractions $(d_{g})_{g\in G}\in A$ satisfying $d=\sum_{g\in G}d_{g}\in A^{\alpha}$, such that the following hold:
@@
This follows from [27, Example 5.10] (see also [1, Theorem 3.10]) and Proposition 6.3. ∎
@@
Let $n\in\mathbb{N}$, let $A_{1},A_{2},\dots,A_{n}$ be unital C*-algebras, and for $m=1,2,\dots,n$ let $e_{m}\in A_{m}$ be a positive contraction and let $\tau_{m}$ be a tracial state on $A_{m}$. Let $A=A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}$ (minimal tensor product), and let
@@
Let $n\in\mathbb{N}$, let $A_{1},A_{2},\dots,A_{n}$ be unital C*-algebras, and for $m=1,2,\dots,n$ let $e_{m}\in A_{m}$ be a positive contraction and let $\tau_{m}$ be a tracial state on $A_{m}$. Let $A=A_{1}\otimes A_{2}\otimes\dots\otimes A_{n}$ (minimal tensor product), and let
@@
Let $F\subseteq A$ and $S\subseteq C(G)$ be finite sets, let $\varepsilon>0$, let $x\in A_{+}$ with $\|x\|=1$, and let $y\in(A^{\alpha})_{+}\setminus\{0\}$. Without loss of generality, we may assume $\|a\|\leq 1$ for all $a\in F$, $\|f\|\leq 1$ for all $f\in S$, and $\varepsilon<1$. We need to find a positive contraction $d\in A^{\alpha}$ and a completely positive contractive map $\psi\colon C(G)\to\overline{dAd}$ such that the following hold:
@@
By [27, Example 5.10], the action $\gamma$ has the weak tracial Rokhlin property. Therefore, by [13, Lemma 3.3], $B\rtimes_{\gamma}S_{2}$ is simple, and hence by Theorem 4.5, $B^{\gamma}$ is simple. It is easy to verify that $A^{\alpha}$ can be identified with $\bigotimes_{m=1}^{\infty}B_{m}^{\gamma}$ (see [37, Lemma 4.5]), hence it is simple. Meanwhile, since $B$ has strict comparison, $B^{\gamma}$ also has it, and thus $A^{\alpha}$ does as well (see [12, Proposition 5.16.1(4)]). Since $A^{\alpha}$ is unital, stably finite, and amenable, $\mathrm{T}(A^{\alpha})$ is a nonempty compact set. We denote $\delta_{2}=\inf_{\tau\in\mathrm{T}(A^{\alpha})}d_{\tau}(y)>0$.
@@
Since the action $\gamma$ has the weak tracial Rokhlin property, by Lemma 7.4, for each $1\leq i\leq N$, set $F_{i}=\{c_{ij}\colon 1\leq j\leq k\}$, we obtain a positive contraction $d^{(i)}\in B^{\gamma}$ and mutually orthogonal positive contractions $(d_{g}^{(i)})_{g\in H_{i}}\in B$ such that the following hold:
@@
In fact, we can use the above technique to construct examples of permutation actions on unital simple tracially $\mathcal{Z}$-stable C*-algebras with few projections. For example, the algebras constructed in [14] are all unital simple tracially $\mathcal{Z}$-stable with few projections, have strict comparison, and stable rank one (hence stably finite). The verification of conditions (3) - (5) can also be performed using the approach in the proof of [37, Theorem 4.8].
Funding
No funding was received.